# Fluctuations of local plastic strain in granular media

### v1

@@ Abstract
We experimentally study the heterogeneity of strain in a granular medium subjected to oscillatory shear in a rotating drum. Two complementary methods are used. The first method relies on optical imaging and grain tracking, allowing us to compute some components of the strain tensor and their variance. The second method, Diffusive Acoustic Wave Spectroscopy (DAWS), provides the variance of strain components within the bulk. Our results show that strain is spatially heterogeneous, with fluctuations about ten times larger than the mean, primarily dominated by variability at the grain scale. We then analyze in detail the strain fluctuations occurring during the forward and backward branches of the shear cycles, as well as the plastic strain at the completion of each cycle. Both methods reveal that each shear cycle consists of two consecutive diffusive-like branches, and that the resulting plastic strain fluctuations grow linearly with shear. We propose a simple framework to rationalize these effects, suggesting that plastic strain fluctuations increase with the diffusion coefficient and decrease with the anticorrelation between forward and backward strain, which is enhanced in denser packings.
@@ I Introduction
This paper has two main objectives. First, it proposes an method for monitoring strain fluctuations at the grain scale in sheared granular media. Second, it offers new insights into the dependence of the intensity of local strain heterogeneity on macroscopic shear.
@@ II Experiments
Two series of experiments were conducted using the same preparation protocol but different measurement techniques. The first method involves analyzing the two-dimensional kinematics by tracking the positions of grains located at the front face of the drum. The second method, DAWS, is a three-dimensional probe of the granular medium and relies on the correlation of ultrasound signals propagating through the water pores.
II.1 Preparation protocol of the initial state
@@
A sketch of the rotating drum is provided in Figure 1 a), b), and c). After introducing the grains (ceramic, diameter $d=1.50\pm 0.08$ mm), the drum is mounted on a rotating stage. It is then filled with water, ensuring the removal of any air bubbles. Each experiment is preceded by a rapid, continuous clockwise rotation of the drum by at least $180^{\circ}$. The preparation protocol then consists of a quasi-static anticlockwise rotation, inducing approximately twenty consecutive avalanches (Figure 1 d)). This quasi-static rotation allows us to measure two characteristic angles, defined as the inclination $\theta$ of the granular free surface relative to the horizontal plane.
@@
Figure 1: Sketch of the experiment. The rotating drum (radius $D/2=100$ mm and thickness $L=20$ mm) is filled with grains and water. (a) The front view shows the grains and the three zones of interest analyzed. (b) The side view for the grain tracking method shows the camera, which rotates with the drum. Only the grains at the front face are imaged. (c) The side view for the DAWS setup illustrates the ultrasound emission and reception transducers. The thick white line represents an arbitrary diffusive path. (d) For both the grain tracking method and DAWS, the preparation protocol begins with the continuous rotation of the drum, producing several consecutive avalanches. Then, the free surface is brought to an angle $+\theta_{m}$, lower than the angle of repose, and subjected to 50 oscillations between $+\theta_{m}$ and $-\theta_{m}$. Black points indicate the angles at which images or ultrasonic signals are recorded. The notations $a_{k}$ and $c_{k}$ are introduced later in the text to define the pairs of images (two granular configurations) between which the strain is analyzed.
@@
The avalanche angle, $\theta_{a}=25.5^{\circ}$, represents the maximum inclination before an avalanche occurs, while $\theta_{r}=22.8^{\circ}$ is the inclination at which the granular medium stops flowing [19]. Consequently, the preparation protocol ensures that the free surface of the granular medium is initially inclined at $\theta\approx\theta_{r}$. At this stage, the drum is rotated so that the free surface is set to an inclination $\theta_{m}<\theta_{r}$ (Fig. 1d)). The system is then subjected to 50 oscillations at $\theta=\pm\theta_{m}$, indexed by $1\leq k\leq 50$. We present results for four different values of $\theta_{m}$: $21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$. For ensemble averaging, the entire preparation protocol and the subsequent fifty shear cycles are repeated 10 times for each $\theta_{m}$.
@@ II.2 Image analysis
The granular medium is characterized by capturing four images per cycle of the grains at the front face of the drum using the camera, as indicated by the black dots in Figure 1d). An example of such an image is shown in Figure 2a). We present data for three different areas of interest, fixed on the side of the drum. Two of these areas are near the free surface (P1111 and P3333), while the third one is located deeper within the packing (P2222), as shown in Figure 2a).
@@ II.2.1 Grain detection and tracking
During the initial oscillations, grain motion can be significant ($>d/2$), leading to inaccurate tracking. To ensure robust analysis, we retain only images where grain tracking can be visually validated. The first cycle in which grains are accurately tracked is $k_{r}=4,3,2,1$ for $\theta_{m}=21.8^{\circ},19.8^{\circ},16.8^{\circ}$, and $10.8^{\circ}$, respectively. This first cycle is marked with a star in Figure 3a).
Figure 2: Typical image taken at the front face of the drum, illustrating examples of grain detection and tracking. a) Each detected grain is arbitrarily colored. The positions ”$P1$”, ”$P2$”, and ”$P3$”, outlined by red squares, correspond to the zones of interest where deformation is computed and compared with the decorrelation of acoustic signals. b) Reconstructed trajectories over $50$ oscillations at position $P1$, located within the black square in (a). c) Zoom on a selected trajectory from (b).
@@ II.2.2 Strain analysis and averaging procedure
The optical imaging method allows us to compute the strain components $\langle\overline{\epsilon_{xx}}\rangle$, $\langle\overline{\epsilon_{yy}}\rangle$, $\langle\overline{\epsilon_{xy}}\rangle$, and $\langle\overline{\epsilon_{v}}\rangle$, while $\langle\overline{\epsilon_{zz}}\rangle=0$ due to the presence of solid walls. In addition to being limited to two dimensions, optical imaging captures the kinematics of grains at the front face of the drum, which may behave differently from those within the bulk. For instance, it is well known that in flowing granular media, solid walls induce layering, density oscillations, and inhomogeneous granular temperature [23].
@@ II.2.3 Free surface
A useful analysis is the detection of the free surface to estimate the variation in the global solid fraction throughout the experiments. We first divide the image into 30 vertical stripes of equal width (4d). Then, we compute the profile of intensity for each stripe and apply a detection threshold to find the position of the free surface $y_{s}$ for each stripe. We calculate its average position throughout the width of the drum $\overline{y_{s}}$ followed by ensemble averaging. The vertical displacement of the free surface from its initial position that was reached at the end of the protocol of preparation is defined as $\langle\overline{-s_{y}}\rangle=\langle\overline{y_{s}(k)}\rangle-\langle\overline{y_{s}(0)}\rangle$. Unlike grain tracking, free surface detection can be performed as early as the first oscillation cycle. For comparison, the first cycles from which grain tracking could be performed are annotated by stars in Figure 3 a).
@@ II.3 Diffusing acoustic wave spectroscopy
It is also desirable to probe the granular medium in the bulk, which is achieved using DAWS. When the wavelength of a propagating wave is on the order of the particle size, multiple scattering occurs due to the numerous scatterers encountered along the wave’s path. DAWS relies on analyzing the variations in the phase variance of the propagating signal along the acoustic path of $n$ scatterers (grains), caused by their relative motion [15, 24].
@@ III.1 Compaction
Figure 3: a) Variation of the free surface position as a function of the number of oscillations. The curves, from lowest to highest, correspond to $\theta_{m}=21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$. The black lines represent the fit from the model proposed by [26]. The fitted parameters are provided in the appendix (see Fig. 10). A star denotes the first images where grain tracking could be performed. b) Profile of the volumetric strain [$a_{k_{r}},a_{50}$] along the vertical axis $\langle\overline{\epsilon_{v}(\theta_{m})}\rangle$, normalized by the mean value $\epsilon^{*}$ over the region $0<|y|<60~{}\text{mm}$. The colors red, blue, green, and pink correspond to $\theta_{m}=21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$, respectively. Experimental data are represented by circles, while the thin lines represent the exponential fit (as described in the text).
As the number of oscillations increases, the granular material undergoes compaction, as evidenced by the variation of $\langle\overline{-s_{y}}\rangle$ shown in Figure 3 a). Higher oscillation angles result in denser packings for a fixed number of oscillations. Additionally, compaction occurs more rapidly at larger oscillation angles. The evolution of packing fraction during shaking is fitted to the standard logarithmic law, as proposed by [26], and is presented in Figure 3 a). The equation used for the fitting and the corresponding parameters are provided in Appendix 10, Figure 10. It can be observed that the quasi-static oscillations employed in this study have a similar effect to vertical mechanical taps [26] in terms of compacting the granular material.
Furthermore, as reported in [26], the compaction is spatially heterogeneous, being more pronounced near the free surface than in the bulk of the granular medium. To capture this spatial variation with sufficient resolution, we apply a specific binning technique along the vertical direction by dividing the drum images into horizontal stripes of thickness $4d$, and compute the strain for each stripe. The profiles of volumetric strain between [$a_{k_{r}},a_{50}$] are shown in Figure 3 b) as a function of depth, along with an exponential fit of the form $\langle\overline{\epsilon_{v}}\rangle\propto\exp\left(-\frac{y}{y^{*}}\right)$. A similar profile is observed for each oscillation angle $\theta_{m}$, with the characteristic cutoff length $y^{*}=11.4\pm 3$ mm.
@@ III.2.1 $\theta\text{-}\epsilon$ cycles
In Figure 4a), we plot $\theta\text{-}\epsilon$ cycles, focusing on the shear strain $\langle\overline{\epsilon_{xy}}\rangle$ at position $P1$ for $\theta_{m}=19.8^{\circ}$. As the granular material compacts, the maximum intra-cycle shear strain $\langle\overline{\epsilon_{xy,m}}\rangle$, defined as the strain between $[c_{k},a_{k+1}]$ and illustrated in Figure 4b), decreases with the number of oscillations. Additionally, between $[a_{k},a_{k+1}]$, each cycle is characterized by finite plastic deformation $\langle\overline{\epsilon_{xy,p}}\rangle$ (see also Figure 4b). After a large number of oscillations, the cycles stabilize and the plastic strain disappears. This stabilization, along with permanent cumulative deformation, is referred to as ”accommodation” in soil mechanics [27].
@@
An example of a deformation map for the interval $[a_{k_{r}},a_{50}]$ is shown in Figure 5 a), where both volumetric and shear strains are presented. The strain distribution is spatially heterogeneous across two length scales. At the grain scale, adjacent Bagi cells can exhibit strain differences by several orders of magnitude, with strains that can even have opposite signs. At the larger scale represented by the characteristic cutoff length $y^{*}$ in Figure 3 b), the strain map shows that that strain is more pronounced near the free surface.
@@
In Figure 5 b), we present the strain and its root-mean-square fluctuation $\delta\epsilon_{p}=\left(\langle\overline{\epsilon_{p}^{2}}\rangle-\langle\overline{\epsilon_{p}}\rangle^{2}\right)^{1/2}$ for the interval $[a_{k_{r}},a_{k}]$, which corresponds to the total plastic deformation accumulated during a given cycle. For both the shear and volumetric components, the root-mean-square fluctuations are approximately ten times greater than the mean, with $\delta\epsilon_{p}\approx\langle\overline{\epsilon_{p}^{2}}\rangle^{1/2}\gg|\langle\overline{\epsilon_{p}}\rangle|$. This is not attributable to the heterogeneity of the mean strain, as an exponential profile leads to a variability of the same order of magnitude as the mean. Furthermore, the fluctuations are not influenced by the reproducibility of $\langle\overline{\epsilon_{p}}\rangle$ across different experiments, as shown by the error bars in Figure 5 b), which are much smaller than the fluctuations themselves. Therefore, the fluctuations are primarily governed by the strain heterogeneity at the grain scale, caused by the erratic motions of the grains. These motions are the dominant source of decorrelation in the scattered waves, which is quantified by $g_{1}$. From this point onward, we will refer to $\delta\epsilon\approx\langle\overline{\epsilon^{2}}\rangle^{1/2}$ as fluctuations at the grain scale.
@@
Figure 5: a) Maps of Bagi cells for plastic volumetric strain (upper panel) and plastic shear strain (lower panel) for $\theta_{m}=19.8^{\circ}$, pair $[a_{k_{r}},a_{50}]$. The upper portion of each panel corresponds to the free surface. b) Plot of $|\langle\epsilon_{v}\rangle|$, $|\langle\epsilon_{xy}\rangle|$, $\langle\epsilon^{2}_{v}\rangle^{1/2}$, $\langle\epsilon^{2}_{xy}\rangle^{1/2}$, and $\delta\epsilon_{v}$, $\delta\epsilon_{xy}$ as a function of the number of oscillations $k$, at position $P1$ for $\theta_{m}=19.8^{\circ}$. Data are presented for pairs $[a_{k_{r}},a_{k}]$. The ordinate is shown on a logarithmic scale for convenient visualization of the data. The error bars on the $|\langle\epsilon_{v}\rangle|$ and $|\langle\epsilon_{xy}\rangle|$ curves represent the standard deviation arising from the variability in the mean deformation between different experiments (10 in total).
@@
Figure 6: a) Quadratic plastic strain components as a function of the number of oscillations. b) Distribution of the relative displacements $\langle\Delta r_{x}\rangle$ and $\langle\Delta r_{y}\rangle$ between cycles $k_{r}$ and $50$. In both panels a) and b), the data correspond to $\theta_{m}=16.8^{\circ}$, position $P1$, and $k_{r}=2$.

### v2

@@ Abstract
We experimentally study the heterogeneity of strain in a granular medium subjected to oscillatory shear in a rotating drum. Two complementary methods are used. The first method relies on optical imaging and grain tracking, allowing us to compute some components of the strain tensor and their variance. The second method, Diffusive Acoustic Wave Spectroscopy (DAWS), provides the quadratic strain within the bulk. Our results show that strain is spatially heterogeneous, with fluctuations about ten times larger than the mean, primarily dominated by variability at the grain scale. We then analyze in detail the strain fluctuations occurring during the forward and backward branches of the shear stress cycles, along with the intra-cycle plastic strain resulting from of each cycle. Both methods reveal that each shear cycle consists of two consecutive diffusive-like branches, and that the resulting plastic strain fluctuations scales with the mean plastic shear strain. We propose that plastic strain fluctuations result from irreversible strain heterogeneity that increases with applied shear—reflected in forward–backward strain anticorrelations—but is constrained by load-controlled induced memory.
@@ I Introduction
This paper has two aims: to introduce a method for monitoring grain-scale strain fluctuations in sheared granular media and to provide new insights into how the magnitude of local strain heterogeneity depends on the applied macroscopic shear.
@@ II Experiments
Two series of experiments were conducted using the same preparation protocol but different measurement techniquesmethods. The first method involves analyzing the two-dimensional kinematics by tracking the positions of grains located at the front face of the drum. The second method, DAWS, is a three-dimensional probe of the granular medium and relies on the correlation of ultrasound signals propagating through the water pores.
II.1 Preparation protocol of the initial state and oscillations
@@
A sketch of the rotating drum is provided in Figure 1 a), b), and c). After introducing the grains (ceramic, diameter $d=1.50\pm 0.08$ mm), the drum is mounted on a rotating stage. It is then filled with water, ensuring the removal of any air bubbles. Each experiment is preceded by a rapid, continuous anticlockwise rotation of the drum by at least $180^{\circ}$. The preparation protocol then consists of a slower clockwise rotation, inducing approximately twenty consecutive avalanches (Figure 1 d)). This quasi-static rotation allows us to measure two characteristic angles, defined as the inclination $\theta$ of the granular free surface relative to the horizontal plane.
@@
Figure 1: Sketch of the experiment. The rotating drum (radius $D/2=100$ mm and thickness $L=20$ mm) is filled with grains and water. (a) The front view shows the grains and the three zones of interest analyzed. (b) The side view for the grain tracking method shows the camera (Basler acA1920), which rotates with the drum. Only the grains at the front face are imaged. (c) The side view for the DAWS setup illustrates the ultrasound emission and reception transducers. The thick white line represents an arbitrary diffusive path. (d) For both the grain tracking method and DAWS, the preparation protocol begins with the continuous rotation of the drum, producing several consecutive avalanches. Then, the free surface is brought to an angle $\theta_{m}$, lower than the angle of repose, and subjected to 50 oscillations between $\theta_{m}$ and $\theta_{m}$. Black points indicate the angles at which images or ultrasonic signals are recorded. The notations $a_{k}$ and $c_{k}$ are introduced later in the text to define the pairs of images (two granular configurations) between which the strain is analyzed.
@@
The avalanche angle, $\theta_{a}=25.5^{\circ}$, represents the maximum inclination before an avalanche occurs, while $\theta_{r}=22.8^{\circ}$ is the inclination at which the granular medium stops flowing [19]. Consequently, the preparation protocol ensures that the free surface of the granular medium is initially inclined at $\theta\approx\theta_{r}$.
At this stage, the drum is rotated so that the free surface is set to an inclination $\theta_{m}<\theta_{r}$ (Fig. 1d)). The system is then subjected to 50 oscillations at $\theta=\pm\theta_{m}$, indexed by $1\leq k\leq 50$. The granular medium is therefore studied in the plastic ”solid” regime. As will be shown later, this corresponds to a typical shear strain of $\epsilon\sim 10^{-3}\rightarrow 10^{-2}$ smaller than the yield strain but much larger than the strain characteristic of the elastic regime where $\epsilon\sim 10^{-6}$. To mitigate inertial effects, the rotation between positions labeled $a_{k}$ and $a_{k+1}$ (Fig. 1d)) is performed in four phases combining constant—and deliberately small—acceleration and deceleration stages: (i) acceleration at $0.2^{\circ}/\mathrm{s}^{2}$, (ii) stationary rotation phase at a constant angular velocity of $0.5^{\circ}/\mathrm{s}$, and (iii) deceleration phase at $0.2^{\circ}/\mathrm{s}^{2}$. Furthermore, after reaching the target angular position (black dots Fig. 1d)), a waiting period of $30~\mathrm{s}$ is imposed to ensure all grain motion fully relaxes before acquisition. Ten consecutive images are taken at each position to allow for averaging and reduce residual noise. We present results for four different values of $\theta_{m}$: $21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$. For ensemble averaging, the entire preparation protocol and the subsequent fifty shear cycles are repeated 10 times for each $\theta_{m}$.
@@ II.2 Image analysis
The granular medium is characterized by capturing four images per cycle. An example of such an image is shown in Figure 2 a). The camera is rigidly mounted on the frame of the rotating stage, so that images are recorded in the reference frame of the camera. We present data for three different areas of interest, fixed on the side of the drum. Two of these areas are near the free surface (P11 and P33), while the third one is located deeper within the packing (P22), as shown in Figure 2a).
@@ II.2.1 Grain detection and tracking
During the initial oscillations, grain motion can be significant (>d/2>d/2), leading to inaccurate tracking. To ensure robust analysis, we retain only images where grain tracking can be visually validated. The first cycle in which grains are accurately tracked is $k_{r}=4,3,2,1$ for $\theta_{m}=21.8^{\circ},19.8^{\circ},16.8^{\circ}$, and $10.8^{\circ}$, respectively. This first cycle is marked with a star in Figure 3a).
Figure 2: a) Typical image taken at the front face of the drum, illustrating examples of grain detection the end of the preparation protocol for $\theta_{m}=19.8^{\circ}$. The black arrow represents the direction of gravity $\vec{g}$ at that angle. Each detected grain is arbitrarily colored. The positions ”P1P1”, ”P2P2”, and ”P3P3”, outlined by red squares, correspond to the zones of interest where deformation is computed and compared with the decorrelation of acoustic signals. b) Reconstructed trajectories after tracking for 50 cycles at $P1$ and oscillation angle $\theta_{m}=19.8^{\circ}$. The field of view corresponds to the black square in a. c) Zoom on a selected trajectory from b). The first point $a_{3}$ is located on the top-right part of the image.
@@ II.2.2 Strain analysis and averaging procedure
The optical imaging method allows us to compute the strain components $\langle\overline{\epsilon_{xx}}\rangle$, $\langle\overline{\epsilon_{yy}}\rangle$, $\langle\overline{\epsilon_{xy}}\rangle$, and $\langle\overline{\epsilon_{v}}\rangle$ (volumetric), while $\langle\overline{\epsilon_{zz}}\rangle=0$ due to the presence of solid walls. In addition to being limited to two dimensions, optical imaging captures the kinematics of grains at the front face of the drum, which may behave differently from those within the bulk. For instance, it is well known that in flowing granular media, solid walls induce layering, density oscillations, and inhomogeneous granular temperature [23].
@@ II.2.3 Free surface
A useful analysis is the detection of the free surface to estimate the variation in the global solid fraction throughout the experiments. We first divide the image into 30 vertical stripes of equal width (4d). Then, we compute the profile of intensity for each stripe and apply a detection threshold to find the position of the free surface $y_{s}$ for each stripe. We calculate its average position throughout the width of the drum $\overline{y_{s}}$ followed by ensemble averaging. The vertical displacement of the free surface from its initial position that was reached at the end of the protocol of preparation is defined as $\langle\overline{-s_{y}}\rangle=\langle\overline{y_{s}(k)}\rangle-\langle\overline{y_{s}(0)}\rangle$. Unlike grain tracking, free surface detection can be performed as early as the first oscillation cycle. For comparison, the first cycles from which grain tracking could be performed are annotated by stars in Figure 3 b).
@@ II.3 Diffusing acoustic wave spectroscopy
It is also possible to probe the granular medium in the bulk, which is achieved using DAWS. When the wavelength of a propagating wave is on the order of the particle size, multiple scattering occurs due to the numerous scatterers encountered along the wave’s path. DAWS relies on analyzing the variations in the phase variance of the propagating signal along the acoustic path of $n$ scatterers (grains), caused by their relative motion [15, 24].
@@ III.1 Compaction
Figure 3: a) Variation of the free surface position as a function of the number of oscillations. The curves, from lowest to highest, correspond to $\theta_{m}=21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$. The black lines represent the fit from the model proposed by [26]. The fitted parameters are provided in the appendix (see Fig. 10). A star denotes the first images where grain tracking could be performed. b) Profile of the volumetric strain $a_{k_{r}},a_{50}$] along the vertical axis $\langle\overline{\epsilon_{v}(\theta_{m})}\rangle$, normalized by the mean value $\epsilon^{*}$ over the region $0<|y|<60~\text{mm}$. The colors red, blue, green, and pink correspond to $\theta_{m}=21.8^{\circ}$, $19.8^{\circ}$, $16.8^{\circ}$, and $10.8^{\circ}$, respectively. Experimental data are represented by circles, while the thin lines represent the exponential fit (as described in the text).
As the number of oscillations increases, the granular material undergoes compaction, as evidenced by the variation of $\langle\overline{-s_{y}}\rangle$ shown in Figure 3 a). Higher oscillation angles result in denser packings for a fixed number of oscillations. Additionally, compaction occurs more rapidly at larger oscillation angles. The evolution of packing fraction during compaction is fitted to the standard logarithmic law, as proposed by [26], and is presented in Figure 3 a). The equation used for the fitting and the corresponding parameters are provided in Appendix 10, Figure 10. It can be observed that the quasi-static oscillations employed in this study produce an effect similar to vertical mechanical taps [26] in terms of compacting the granular material. However, this analogy should be interpreted with caution, as the underlying mechanisms may differ: while classical tapping induces compaction through inertial driving, the present oscillations primarily promote compaction via quasi-static shearing.
Furthermore, as shown by [26], the compaction is spatially heterogeneous, being more pronounced near the free surface than in the bulk of the granular medium. To capture this spatial variation with sufficient resolution, we apply a specific binning technique along the vertical direction by dividing the drum images into horizontal stripes of thickness $4d$, and compute the strain for each stripe. The profiles of volumetric strain between $a_{k_{r}},a_{50}$] are shown in Figure 3 b) as a function of depth, along with an exponential fit of the form $\langle\overline{\epsilon_{v}}\rangle\propto\exp\left(-\frac{y}{y^{*}}\right)$. A similar profile is observed for each oscillation angle $\theta_{m}$, with the characteristic cutoff length $y^{*}=11.4\pm 3$ $m$.
@@ III.2.1 $\theta\text{-}\epsilon$ cycles
In Figure 4a), we plot $\theta\text{-}\epsilon$ cycles, focusing on the shear strain $\langle\overline{\epsilon_{xy}}\rangle$ at position $P1$ for $\theta_{m}=19.8^{\circ}$. As the granular material compacts, the maximum intra-cycle shear strain $\langle\overline{\epsilon_{xy,m}}\rangle$, defined as the strain between $a_{k},c_{k}]$ and illustrated in Figure 4b), decreases with the number of oscillations.
To estimate the packing fraction $\phi$ of the material located in the zones $P1$ and $P3$, we use the relation $\phi\approx\phi_{0}\left(1+\langle\overline{s_{y}}\rangle/h\right)$, where $\langle\overline{s_{y}}\rangle$ is the position of the free surface, as shown in Figure 3 a). The difference in packing fraction between two systems sheared at different $\theta_{m}$ is given by $\Delta\phi/\phi_{0}\approx\langle\overline{s_{y}}\rangle/h\ll\Delta\theta_{m}/\theta_{m}$. Moreover, at a fixed $\theta_{m}$, the relative variation of packing fraction during compaction is small, on the order of $\sim 10^{-2}$. Therefore, $\theta_{m}$ determines the constant maximum stress applied at each cycle, independent of $\phi$. However, due to the repeated oscillations, $\theta_{m}$ do modify the packing fraction by also controlling the intensity of ”mechanical taps”, as discussed above. Therefore, the constitutive relation describing the cycles can be expressed as $\langle\overline{\epsilon_{xy,m}}\rangle(\theta_{m},\phi(\theta_{m},k))$, increasing function of $\theta_{m}$ at fixed $\phi$ and decreasing function of $\phi$. In turn, $\phi$ is an increasing function of both $k$ and $\theta_{m}$.
Additionally, between $a_{k},a_{k+1}]$, each cycle is characterized by finite plastic deformation $\langle\overline{\epsilon_{xy,p}}\rangle$ (see also Figure 4b). After a large number of oscillations, the cycles stabilize and the plastic strain disappears. This stabilization, along with permanent cumulative deformation, is referred to as ”accommodation” in soil mechanics [27].
@@
An example of a deformation map for the interval $a_{k_{r}},a_{50}]$ is shown in Figure 5 a), where both volumetric and shear cumulative plastic strains are presented. The strain distribution is spatially heterogeneous across two length scales. At the grain scale, adjacent Bagi cells can exhibit strain differences by several orders of magnitude, with strains that can even have opposite signs. At the larger scale represented by the characteristic cutoff length $y^{*}$ in Figure 3 b), the strain map shows that that strain is more pronounced near the free surface.
@@
In Figure 5 b), we present the cumulative strain and its root-mean-square fluctuation $\delta\epsilon_{p}=\left(\langle\overline{\epsilon_{p}^{2}}\rangle-\langle\overline{\epsilon_{p}}\rangle^{2}\right)^{1/2}$ for the interval $a_{k_{r}},a_{k}]$, which corresponds to the total plastic deformation accumulated during a given cycle. For both the shear and volumetric components, the root-mean-square fluctuations are approximately five to ten times greater than the mean, with $\delta\epsilon_{p}\approx\langle\overline{\epsilon_{p}^{2}}\rangle^{1/2}\gg|\langle\overline{\epsilon_{p}}\rangle|$. This is not attributable to the heterogeneity of the mean strain, as an exponential profile leads to a variability of the same order of magnitude as the mean. Furthermore, the fluctuations are not influenced by the reproducibility of $\langle\overline{\epsilon_{p}}\rangle$ across different experiments, as shown by the error bars in Figure 5 b), which are much smaller than the fluctuations themselves. Therefore, the fluctuations are primarily governed by the strain heterogeneity at the grain scale, caused by the erratic motions of the grains. These motions are the dominant source of decorrelation in the scattered waves, which is quantified by $g_{1}$. From this point onward, we will refer to $\delta\epsilon\approx\langle\overline{\epsilon^{2}}\rangle^{1/2}$ as fluctuations at the grain scale.
@@
Figure 5: a) Maps of Bagi cells for cumulative plastic volumetric strain (upper panel) and plastic shear strain (lower panel) for $\theta_{m}=19.8^{\circ}$, pair $a_{k_{r}},a_{50}]$. The upper portion of each panel corresponds to the free surface. b) Plot of cumulative $|\langle\epsilon_{v,p}\rangle|$, $|\langle\epsilon_{xy,p}\rangle|$, $\langle\epsilon^{2}_{v,p}\rangle^{1/2}$, $\langle\epsilon^{2}_{xy,p}\rangle^{1/2}$, and $\delta\epsilon_{v,p}$, $\delta\epsilon_{xy,p}$ as a function of the number of oscillations $k$, at position $P1$ for $\theta_{m}=19.8^{\circ}$. Data are presented for pairs $a_{k_{r}},a_{k}]$. The ordinate is shown on a logarithmic scale for convenient visualization of the data. The error bars on the $|\langle\epsilon_{v}\rangle|$ and $|\langle\epsilon_{xy}\rangle|$ curves represent the standard deviation arising from the variability in the mean deformation between different experiments (10 in total).
@@
Figure 6: a) Cumulative quadratic plastic strain components as a function of the number of oscillations. b) Distribution of the relative displacements $\langle\Delta r_{x}\rangle$ and $\langle\Delta r_{y}\rangle$ between cycles $k_{r}$ and $50$. In both panels a) and b), the data correspond to $\theta_{m}=16.8^{\circ}$, position $P1$, and $k_{r}=2$.